OpenStudy (anonymous):

Part 1: Use complete sentences to describe a real-world scenario that could be represented by the inequality 4x + 3y greater than or equal to 36. Part 2: Choose one ordered pair that is a solution to the given inequality and explain what that ordered pair means in the context of your real-world scenario.

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

beers cost \(\$3\) and shots are \(\$4\) if i spent over \($36\) what possible combinations of beers and shots could i have had?

OpenStudy (anonymous):

thanks, help on part 2?

OpenStudy (anonymous):

you can pick pretty much anything you like for example if i had 5 shots and 6 beers \((5,6)\) then the total i spent was \[4\times 5+6\times 3=20+18=38\]

OpenStudy (anonymous):

ok